If the proportion of american adults who believe that America is ready for a woman president is .70 what are the mean and standard deviation for the number of people who believes that america is ready for a sample size of 1000?

find the standard deviation and mean.
Answer #1

Solution

Given that,

p = 0.70

1 - p = 1-0.70=0.30

n = 100

= p =0.70

standard deviation and mean. (proportion)

=  [p( 1 - p ) / n] = [(0.70*0.30) / 100]=0.0458

=0.046